OpenStudy (anonymous):

I'll do anything for the answer simplify the radical, square root150x^3k^4

OpenStudy (anonymous):

\[\large \sqrt{150x^3k^4}\] \[150=25\times 6\] \[\sqrt{150}=\sqrt{25\times 6}=\sqrt{25}\sqrt{6}=5\sqrt{6}\] \[\sqrt{x^3}=x\sqrt{x}\]\[\sqrt{k^4}=k^2\]

OpenStudy (anonymous):

put it all together and get \[\large \sqrt{150x^3k^4}=5xk^2\sqrt{6x}\]

OpenStudy (anonymous):

Thank you so much :) <3 <3

OpenStudy (anonymous):

yw